Let [x] denotes greatest integer Then
equal to




Possible Answers























Selected Possible Answer



x/2







x/3







0







x/4





Grade:12

2 Answers

satish prakash garg
14 years ago

x/3 is ans

[x]=x-{x}

and 

sum of sq. of n no. are  n*(n+1)*(2n+1)/6

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Geetika Chouhan,

Ans:-Now the given function is a box function. So [x]=x-{x}

where {x}=fractional part of x whose value lie between 0 and 1

So writing all the box functions like this we get

the series (1²x+2²x+...........n²x) -K where K is a constant whose max value is n

Hence the sum is =  xn(n+1)(2n+1)/6   -n

dividing it by n^3 we get

Sum=x(1+1/n)(2+1/n)/6        -1/n²

Taking n→∞ We get

Sum=x/3(ans)

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Geetika Chouhan !!!

 


Regards,

Askiitians Experts

EXPERT NAME